# Multi-Agent Trajectory Prediction with Difficulty-Guided Feature Enhancement Network

### v1

@@ Abstract
Autonomous driving, Trajectory Prediction, Behavioral Heterogeneity, Feature Enhancement Network.
@@ I Introduction
Currently, mainstream prediction model architectures can be roughly divided into three steps: feature encoding, feature interaction, and feature decoding. Additionally, some methods improve the accuracy and robustness of prediction models through feature enhancement. These methods involve intermediate modeling of future intentions or trajectories and interact and fuse these with the original encoded features. Although these methods significantly improve model performance compared to mainstream model architectures, they fail to consider the inherent heterogeneity in prediction difficulty among agents, i.e., the natural differences in prediction difficulty for different agents within the scene.
@@
To address potential hazards in driving scenarios, during the decision-making and planning stages of autonomous driving, it is common to consider the heterogeneity of traffic participants’ behaviors [1, 2, 3, 4]. This is primarily manifested in driver social behaviors, intent uncertainty, and scenarios characterized by long-tail distributions [5, 6, 1]. They categorize surrounding agents from different measurement perspectives to enhance the safety and comfort of decision-making and planning tasks. Recently, some studies have incorporated long-tail detection into trajectory prediction tasks [4, 7]. Specifically, they use methods such as contrastive learning or clustering to measure similarity, aiming to reduce the distance in feature space between samples that share similar scene characteristics. Compared to the difficulty of driving scenarios, our method explores the influence of prediction difficulty of different agents within a single scene on the network.
In real driving scenarios, human drivers often predict the future behaviors of most agents subconsciously, relying on intuition. However, when faced with agents that may pose risks and exhibit high prediction difficulty, human drivers tend to carefully consider their behavior and motivations [8]. Inspired by human drivers, we aim to investigate the effectiveness of prediction methods that follow the principle of easy-to-hard in multi-agent trajectory prediction tasks. To mimic the prediction approach of human drivers, our first step is to initially predict the potential intentions and future trajectories of all agents in the scene based on historical trajectories and map information. This step resembles the rough subconscious prediction of human drivers. Then, the prediction difficulty differentiation of agents is performed based on the difficulty coefficient modeling. Additionally, based on difficulty differentiation, we use reliable future trajectories as prior information and utilize feature interaction and fusion for feature enhancement. The significance of this step is that the future behaviors of easily predictable agents impose motion constraints on those agents with higher difficulty. Finally, to better capture local motion features and global relationships, we combine two approaches for scene representation.
As illustrated in the Fig. 1, traditional trajectory prediction methods can be broadly summarized into two stages: Encoding&Interaction and Final Prediction. Initially, contextual encoding is employed to extract scene features from various input information. These features are processed through an interaction module, which aids the model in understanding and capturing the interactions and relationships between different entities (such as vehicles, lane markings, traffic lights, etc.). Finally, these features are passed to a decoding module to obtain the final predicted trajectories and their corresponding probability distributions. The primary difference in our pipeline is the addition of an intermediate module that generates reliable future trajectories, thereby better prediction results can be achieved for agents with higher prediction difficulty.
@@
  • We introduce a novel network architecture that combines the strengths of two scene representations and leverages the differences in prediction difficulty among agents for multi-agent trajectory prediction, thereby improving overall prediction accuracy.
  • We propose a difficulty-guided decoder that adjusts the flow of future trajectories based on prediction difficulty, ensuring that reliable trajectories are utilized in subsequent prediction steps.
  • Inspired by human drivers, our method follows the principle of easy-to-hard. By initially predicting the easier-to-predict agents and using their reliable future trajectories for feature interaction and fusion, we enhance the prediction results for agents with higher prediction difficulty.
The rest of this paper is organized as follows: We briefly review the related work in Section II. We present the technical details of our proposed method in Section III. Then, extensive experiments and analysis are presented in Section IV. Finally, we conclude the paper in Section V.
@@ II Related Work
Heterogeneity in Trajectory Prediction. The inherent heterogeneity in driving trajectories can provide rich additional information for decision making tasks, leading to safer and more flexible decision outcomes. Recent research in trajectory prediction tasks has taken such heterogeneity into account to provide personalised predictions. For example, Chen et al. [9] proposed a distribution discrimination method to predict personalised motion patterns by distinguishing latent distributions. Xing et al. [10] introduced a joint time-series modelling approach that considers different driving styles to predict the trajectory of lead vehicles. In contrast to the emphasis on heterogeneity in the aforementioned works, our research focuses primarily on the prediction difficulty of the agents. For example, Makansi et al. [4] addressed the problem of long-tailed data distribution by exploiting the effects of feature embeddings. The use of contrastive learning in feature embeddings aggregates rare and difficult samples, contributing to improved final prediction performance. Qian et al. [11] introduced a teacher learning strategy based on the difficulty of scene prediction. They predefined the prediction difficulty of the scene and trained the proposed model by first training on easier samples and then on more difficult samples. However, both methods did not consider potential differences in difficulty between agents in multi-agent prediction scenarios. In this work, we designed a difficulty estimation module to estimate potential prediction difficulties in multi-agent prediction scenarios. The future trajectories of agents with lower prediction difficulties are integrated into the scene to enhance information.
@@
Future Feature Enhancement. Compared to conventional pipelines, methods that enhance features using future information allow networks to incorporate future constraints and interactions. These methods can be broadly categorized into two approaches: intention and future trajectories. TNT [12] and PECNet [13] model vehicle intention as latent variables and generate different prediction trajectories conditioned on vehicle intentions, forming predicted trajectories under certain constraint conditions. GANet [14] and ADAPT [15] integrate and fuse vehicle intentions for future feature enhancement through feature interaction. Similarly, Prophnet [16] and FFINet [17] model intermediate latent variables in the form of trajectories, where the output of intermediate decoders is one or multiple trajectories. Additionally, there is a method that models future trajectories through planner sampling, followed by feature enhancement [18, 19].
@@
Most relevant to our work is M2I [20], which exploits potential game relationships between vehicles. It uses a Relation Predictor and a Marginal Trajectory Predictor to input the future trajectories of the leader in the game relationship as prior information into the Conditional Trajectory Predictor. However, we find that this method is highly dependent on the accuracy of the inferred game relationships and future trajectories. Accurately identifying game relationships between two vehicles in long-term predictions proves to be very challenging. Therefore, we use the more easily extractable prediction difficulty as a basis for personalized classification. Our approach is an end-to-end inference model, abandoning the multi-model inference method used by M2I.
Scene Representation. Inspired by the application of convolutional neural networks in image processing, early trajectory prediction methods used rasterized images as the main form of scene representation and then extracted 2D image features through convolution [21]. However, this approach required extensive image rendering processing and suffered from occlusion issues between trajectories and map scenes. To address this, Vectornet [22] introduced a unified scene representation by vectorizing trajectories and map scenes, and utilized MLP to extract scene information.
@@
Broadly, vectorization can be categorized into two types: scene-centric and agent-centric. Since the pioneer of vectorization, Vectornet, was designed for single-agent trajectory prediction tasks, it naturally took the last frame coordinates of the vehicle to be predicted as the scene center, and expressed the trajectories and map vector coordinates of all other agents around the scene center. The benefit of this approach is that it enables the preservation of global relative pose features of various objects in the scene during feature extraction, and facilitates the embedding of some prior information or trajectory anchors [16, 12].
@@
However, due to the demands of multi-agent trajectory prediction tasks, there has been a shift towards agent-centric scene representation methods. Building upon agent-centric approaches, some methods aim to preserve as much information as possible about the mutual spatial relationships of various objects in the scene [23, 24, 25, 26]. To achieve this, they incorporate local pairwise-relative pose features and extract and fuse the relative pose information through networks such as GNNs [27] and Transformers [28]. However, such methods suffer from the limitation that historical and predicted trajectories only retain local motion characteristics, making it inconvenient to embed anchor and prior information. Therefore, we aim to preserve the advantages of each method by integrating the two approaches.
@@ III-A Overview
Recent works [11, 10] have been dedicated to associating similar scenes through contrastive learning or clustering, linking scenes based on their similarity, and enhancing features during inference according to the degree of scene similarity. We propose DGFNet, a network designed for multi-agent trajectory prediction using a Difficulty-Guided Feature Enhancement module for asynchronous interaction between historical trajectories and reliable future trajectories. As shown in Fig. 3, the architecture of DGFNet can be divided into three components: spatio-temporal feature encoding, feature interaction, and trajectory decoding.
@@ III-B Problem formulation
Building on the majority of previous work, such as [29, 22, 12, 25, 1], we assume that upstream perceptual tasks can provide high-quality 2D trajectory tracking data for agents in a coordinate system. At the same time, localisation and mapping tasks can provide accurate self-vehicle positioning data and comprehensive HD map data for the urban environment. In other words, for $j$ agents within a given scene, we obtain $x$ and $y$ positions, denoted as $X^{0:j}_{-t_{h}:0}$, corresponding to the time stamp horizon $\left\{-t_{h},...,0,1,...,t_{f}\right\}$. The trajectory prediction task consists of predicting the future trajectories $Y^{0:j}_{1:t_{f}}$ of the agents by using the HD map information $M$ (including road coordinates, topological links, traffic lights and other relevant information) within the given scene and the historical trajectories output by the tracking task.
In accordance with the encoding methodology employed by VectorNet and LaneGCN, we represent the trajectory information of agents and lane details in vectorized form. To elaborate further, for a given agent $i$ within the scenario, its trajectory information, denoted as $X^{i}$, is represented as a matrix $X^{i}_{-t_{h}:0}$ comprising a spatio-temporal sequence$\left\{s^{i}_{-t_{h}},s^{i}_{-t_{h}+1},...,s^{i}_{0}\right\}$ over the past $t_{h}$ time step. Similarly, map information is also segmented into predefined sequence vectors to capture various scene details. Analogous to trajectory vectors, map information is represented as $M_{1:N}=\left\{s_{1},s_{2},...,s_{N}\right\}$, where $N$ denotes the total vector length. Within each lane vector $s_{i}$, there is an inclusion of lane slice coordinates, lane type (e.g., straight or left-turn lane), and map lane details such as traffic signals. This encapsulates diverse map information pertaining to each lane.
It should be noted that this method involves two types of scenario representations. As shown in Fig. 2, the scene-centric representation involves only one coordinate system, where all agents and lane markings are expressed within this coordinate system. For simplicity, we only use uppercase notation to denote the trajectory and map tensor as $X$ and $M$, respectively. In the agent-centric representation, multiple coordinate systems exist. Initially, all agents and lane markings are rotated using matrix rotation so that their orientations align with the x-axis. Subsequently, we use the original coordinates $p$ and velocity orientation $v$ to calculate the relative position distance $d_{i\rightarrow j}$ and the angular difference $\alpha_{i\rightarrow j}$, $\beta_{i\rightarrow j}$. We denote the agent-centric trajectory and map tensor as $A$ and $L$, respectively. Pairwise-relative pose is represented as $RPE=\left\{||d_{i\rightarrow j}||,\sin(\alpha_{i\rightarrow j}),\cos(\alpha_{i\rightarrow j}),\sin(\beta_{i\rightarrow j}),\cos(\beta_{i\rightarrow j})\right\}$.
@@ III-C Spatio-temporal feature extraction
$Z_{i}=\mathbf{Conv1d}(\mathbf{Res1d}(X_{i}))\quad\text{for }i=0,1,\dots,n_{\text{fpn}}-1$ (1)
$\hat{X}_{i}=\mathbf{Inter}(Z_{i+1},\text{scale\_factor}=2)+Z_{i}$ (2)
after Res1d and Conv1d encoding, feature alignment is performed by scaling and interpolation operations. For the agent-centric trajectory tensor, the same feature extraction is used to obtain $\hat{A}$.
@@
${H}_{1}$ $=\mathbf{PAB}_{1}\left(\mathbf{ReLU}(\mathbf{LayerNorm}(M{W}_{p}+{b}_{p}))\right)$ (3)
$\hat{M}$ $=\mathbf{PAB}_{2}({H}_{1})$ (4)
where $M\in\mathbb{R}^{N_{\text{lane}}\times 10\times C_{\text{in}}}$ are the input lane features, ${W}_{p}\in\mathbb{R}^{C_{\text{in}}\times h}$ and $\mathbf{b}_{p}\in\mathbb{R}^{h}$ are the weights and biases for the projection layer, and PAB represents the PointAggregateBlock.
Similarly, we can obtain the agent-centric map feature $\hat{L}$. For $RPE$ we use MLP for feature extraction to get $\hat{RPE}$.
@@ III-D Feature interaction
We first introduce a generic feature interaction block. In contrast to the general method, which uses simple attention operations for feature interaction, we believe that multi-head attention mechanisms are often better suited to handle complex input sequences and capture a wider range of hierarchical and diverse information. We opt for Multi-Head Attention Blocks (MHAB) instead of the previously used simple attention operations, as used for the features [31]. Specifically, we use self-attention encoders and feedforward networks (FFN) for self-interaction and cross-attention encoders and FFN for cross-interaction:
@@
$\hat{X}^{(l)}=\mathbf{MHA}_{MA}(\hat{M}^{(l-1)},\hat{X}^{(l-1)})$ (10)
$\hat{M}^{(l)}=\mathbf{MHA}_{MM}(\hat{M}^{(l-1)})$ (11)
$\hat{M}^{(l)}=\mathbf{MHA}_{AM}(\hat{X}^{(l)},\hat{M}^{(l-1)})$ (12)
$\hat{X}^{(l)}=\mathbf{MHA}_{AA}(\hat{X}^{(l-1)})$ (13)
Here, $l$ denotes the current layer index, and $\hat{X}^{(0)}=\hat{X}$ and $\hat{M}^{(0)}=\hat{M}$ are the initial inputs.
@@
$\hat{A}^{(k)}=\mathbf{MHA}_{edge}(\hat{A}^{(k-1)},\hat{L}^{(k-1)},\hat{RPE}^{(k-1)})$ (14)
@@ III-E Difficulty-guided decoder
To capture features of plausible future trajectories in the scene, we introduce a Difficulty-Guided Decoder to obtain the future trajectories of agents that are relatively easier to predict. Firstly, we decode the actor features $\hat{A}^{(k)}$ in Agent-centric, which have undergone multiple layers of feature interactions. Each agent receives six predicted trajectories. Following the ADAPT [15] trajectory decoder to get higher quality decoded trajectories by predicting endpoints for refinement. The decoding process can be expressed as follows:
$\hat{E}=\mathbf{MLP}_{\text{end}}(\hat{X}^{(l)})+\mathbf{MLP}_{\text{refine}}(\text{Cat}[\hat{X}^{(l)},\hat{E}])$ (15)
$\hat{P}=\text{Cat}[\mathbf{MLP}_{\text{traj}}(\text{Cat}[\hat{X}^{(l)},\hat{E}]),\hat{E}]$ (16)
Where endpoints $\hat{E}$ are predicted and refined by concatenating with the agent features $\hat{X}^{(l)}$, then the trajectories $\hat{P}$ are predicted and merged with the refined endpoints.
@@
0:  Predicted trajectories $\hat{P}$, Predicted endpoints $\hat{E}$, threshold $\tau$
0:  Masked trajectories $\hat{P}_{masked}$
1:  $\hat{P}_{masked}=set()$
2:  $\bar{E}\leftarrow\hat{E}.mean()$
@@
4:     $D=\|\hat{E}-\bar{E}\|_{2}$
5:     $\bar{D}\leftarrow D.mean()$
6:     if $\bar{D}\leq\tau$ then
7:        $\hat{P}_{masked}$.add($\hat{P}_{l}$)
@@
9:  return  $\hat{P}_{masked}$
The Argoverse 1 validation and train sets show that the predicted trajectories of most of the vehicles in the scenarios exhibit a high degree of concentration, with more than 90% of the samples of straight ahead trajectories in each case [32]. This suggests that the motion patterns of most agents can be easily captured, reflecting real-world traffic scenarios. In order to prevent higher difficulty prediction trajectories from entering the subsequent modules and generating cumulative errors, we introduced a difficulty masker to filter the initial prediction results. ScenceTransformer[33], inspired by recent approaches to language modeling, has pioneered the idea of using a masking strategy as a query to the model, enabling it to invoke a single model to predict agent behavior in multiple ways (marginal or joint prediction). This mechanism in this paper is intended to mask out the initial prediction trajectories of the more difficult agents, thus providing better control over the prediction accuracy. This process can be formalized as Algo. 1.

### v2

@@ Abstract
Autonomous Vehicle Navigation; Deep Learning Methods; Representation Learning.
@@ I Introduction
Currently, mainstream prediction model architectures can be roughly divided into three steps: feature encoding, feature interaction, and feature decoding. Additionally, some methods improve the accuracy and robustness of prediction models through feature enhancement [1, 2, 3, 4, 5, 6, 7, 8, 9]. These methods involve intermediate modeling of future intentions or trajectories and interact and fuse these with the original encoded features. Although these methods significantly improve model performance compared to mainstream model architectures, they neglect the inherent heterogeneity in prediction difficulty among agents, i.e., the natural differences in prediction difficulty for different agents within the scene. As illustrated in the Fig. 1, The primary difference in our pipeline is the addition of an intermediate module that generates reliable future trajectories, thereby better prediction results can be achieved for agents with higher prediction difficulty.
In real driving scenarios, human drivers often predict the future behaviors of most agents subconsciously, relying on intuition. However, when faced with agents that may pose risks and exhibit high prediction difficulty, human drivers tend to carefully consider their behavior and motivations [10]. Inspired by human drivers, we aim to investigate the effectiveness of prediction methods that follow the principle of easy-to-hard in multi-agent trajectory prediction tasks.
@@
1) We introduce a novel network architecture that combines the strengths of two scene representations and leverages the differences in prediction difficulty among agents for multi-agent trajectory prediction, thereby improving overall prediction accuracy.
2) Inspired by human drivers, our method follows the principle of easy-to-hard. By initially predicting the easier-to-predict agents and using their reliable future trajectories for feature interaction and fusion, we enhance the prediction results for agents with higher prediction difficulty.
@@ II Related Work
Future Feature Enhancement. Compared to conventional pipelines, methods that enhance features using future information allow networks to incorporate future constraints and interactions. These methods can be broadly categorized into two approaches: intention and future trajectories. TNT [1] and PECNet [2] model vehicle intention as latent variables and generate different prediction trajectories conditioned on vehicle intentions, forming predicted trajectories under certain constraint conditions. GANet [3] and ADAPT [4] integrate and fuse vehicle intentions for future feature enhancement through feature interaction. Similarly, Prophnet [5] and FFINet [6] model intermediate latent variables in the form of trajectories, where the output of intermediate decoders is one or multiple trajectories. It should be noted that the first prediction results of both Prophnet and FFINet are obtained by decoding only the historical trajectory features, which leads to a large error in both first prediction results. Additionally, there is a method that models future trajectories through planner sampling, followed by feature enhancement [8, 9].
@@
Most relevant to our work is M2I [7], which exploits potential game relationships between vehicles. It uses a Relation Predictor and a Marginal Trajectory Predictor to input the future trajectories of the leader in the game relationship as prior information into the Conditional Trajectory Predictor. However, we find that this method is highly dependent on the accuracy of the inferred game relationships and future trajectories. Therefore, we use the more easily extractable prediction difficulty as a basis for personalized classification.
@@
Scene Representation. Broadly, vectorization can be categorized into two types: scene-centric and agent-centric. Since the pioneer of vectorization, Vectornet, was designed for single-agent trajectory prediction tasks, it naturally took the last frame coordinates of the vehicle to be predicted as the scene center, and expressed the trajectories and map vector coordinates of all other agents around the scene center. The benefit of this approach is that it enables the preservation of global relative pose features of various objects in the scene during feature extraction, and facilitates the embedding of some prior information or trajectory anchors [5, 1].
@@
However, due to the demands of multi-agent trajectory prediction tasks, there has been a shift towards agent-centric scene representation methods. Building upon agent-centric approaches, some methods aim to preserve as much information as possible about the mutual spatial relationships of various objects in the scene [11, 12, 13, 14, 15]. To achieve this, they incorporate local pairwise-relative pose features and extract and fuse the relative pose features through networks. However, such methods suffer from the limitation that historical and predicted trajectories only retain local motion characteristics, making it inconvenient to embed anchor and prior information. Therefore, we aim to preserve the advantages of each method by integrating the two approaches.
@@ III-A Overview
We propose DGFNet, a network designed for multi-agent trajectory prediction using a Difficulty-Guided Feature Enhancement module for asynchronous interaction between historical trajectories and reliable future trajectories. As shown in Fig. 3, the architecture of DGFNet can be divided into three components: spatio-temporal feature encoding, feature interaction, and trajectory decoding.
@@ III-B Problem formulation
Building on the majority of previous work, such as [16, 17, 1, 13, 18], we assume that upstream perceptual tasks can provide high-quality 2D trajectory tracking data for agents in a coordinate system. At the same time, localisation and mapping tasks can provide accurate self-vehicle positioning data and comprehensive HD map data for the urban environment. In other words, for $j$ agents within a given scene, we obtain $x$ and $y$ positions, denoted as $\mathcal{X}^{0:j}_{-t_{h}:0}$, corresponding to the time stamp horizon $\left\{-t_{h},...,0,1,...,t_{f}\right\}$. The trajectory prediction task consists of predicting the future trajectories $\mathcal{Y}^{0:j}_{1:t_{f}}$ of the agents by using the HD map information $\mathcal{M}$ (including road coordinates, topological links, traffic lights ) within the given scene and the historical trajectories output by the tracking task.
We represent the trajectory information of agents and lane details in vectorized form. To elaborate further, for a given agent $i$ within the scenario, its trajectory information, denoted as $\mathcal{X}^{i}$, is represented as a matrix $\mathcal{X}^{i}_{-t_{h}:0}$ comprising a spatio-temporal sequence$\left\{s^{i}_{-t_{h}},s^{i}_{-t_{h}+1},...,s^{i}_{0}\right\}$ over the past $t_{h}$ time step. Similarly, map information is also segmented into predefined sequence vectors to capture various scene details. Analogous to trajectory vectors, map information is represented as $\mathcal{M}_{1:N}=\left\{s_{1},s_{2},...,s_{N}\right\}$, where $N$ denotes the total vector length. Within each lane vector $s_{i}$, there is an inclusion of lane slice coordinates, lane type (e.g., straight or left-turn lane), and map lane details such as traffic signals. This encapsulates diverse map information pertaining to each lane.
It should be noted that this method involves two types of scenario representations. As shown in Fig. 2, the scene-centric representation involves only one coordinate system, where all agents and lane markings are expressed within this coordinate system. For simplicity, we only use uppercase notation to denote the trajectory and map tensor as $\mathcal{X}$ and $\mathcal{M}$, respectively. In the agent-centric representation, multiple coordinate systems exist. Initially, all agents and lane markings are rotated using matrix rotation so that their orientations align with the x-axis. Subsequently, we use the original coordinates $p$ and velocity orientation $v$ to calculate the relative position distance $d_{i\rightarrow j}$ and the angular difference $\alpha_{i\rightarrow j}$, $\beta_{i\rightarrow j}$. We denote the agent-centric trajectory and map tensor as $\mathcal{A}$ and $\mathcal{L}$, respectively. Pairwise-relative pose is represented as $\mathcal{RPE}=\left\{||d_{i\rightarrow j}||,\sin(\alpha_{i\rightarrow j}),\cos(\alpha_{i\rightarrow j}),\sin(\beta_{i\rightarrow j}),\cos(\beta_{i\rightarrow j})\right\}$.
@@ III-C Spatio-temporal feature extraction
$\mathcal{Z}_{i}=\mathbf{Conv1d}(\mathbf{Res1d}(\mathcal{X}_{i}))\quad\text{for }i=0,1,\dots,n_{\text{fpn}}-1$ (1)
$\hat{\mathcal{X}}_{i}=\mathbf{Inter}(\mathcal{Z}_{i+1},\text{scale\_factor}=2)+\mathcal{Z}_{i}$ (2)
after Res1d and Conv1d encoding, feature alignment is performed by scaling and interpolation operations. For the agent-centric trajectory tensor, the same feature extraction is used to obtain $\hat{\mathcal{A}}$.
@@
${\mathcal{H}}_{1}$ $=\mathbf{PAB}_{1}\left(\mathbf{ReLU}(\mathbf{LayerNorm}(\mathcal{M}{W}_{p}+{b}_{p}))\right)$ (3)
$\hat{\mathcal{M}}$ $=\mathbf{PAB}_{2}({\mathcal{H}}_{1})$ (4)
where $\mathcal{M}\in\mathbb{R}^{N_{\text{lane}}\times 10\times C_{\text{in}}}$ are the input lane features, ${W}_{p}\in\mathbb{R}^{C_{\text{in}}\times h}$ and $\mathbf{b}_{p}\in\mathbb{R}^{h}$ are the weights and biases for the projection layer, and PAB represents the PointAggregateBlock.
Similarly, we can obtain the agent-centric map feature $\hat{\mathcal{L}}$. For $\mathcal{RPE}$ we use MLPs for feature extraction to get $\hat{\mathcal{RPE}}$.
@@ III-D Feature interaction
We first introduce a generic feature interaction block. In contrast to the general method, which uses simple attention operations for feature interaction, we believe that multi-head attention mechanisms are often better suited to handle complex input sequences and capture a wider range of hierarchical and diverse information. We opt for Multi-Head Attention Blocks (MHAB) instead of the previously used simple attention operations, as used for the features [20].
@@
$\hat{\mathcal{X}}^{(l)}=\mathbf{MHA}_{MA}(\hat{\mathcal{M}}^{(l-1)},\hat{\mathcal{X}}^{(l-1)})$ (10)
^$\hat{\mathcal{M}}^{(l)}=\mathbf{MHA}_{MM}(\hat{\mathcal{M}}^{(l-1)})$ (11)
^$\hat{\mathcal{M}}^{(l)}=\mathbf{MHA}_{AM}(\hat{\mathcal{X}}^{(l)},\hat{\mathcal{M}}^{(l-1)})$ (12)
$\hat{\mathcal{X}}^{(l)}=\mathbf{MHA}_{AA}(\hat{\mathcal{X}}^{(l-1)})$ (13)
here, $l$ denotes the current layer index, and $\hat{\mathcal{X}}^{(0)}=\hat{\mathcal{X}}$ and ^$\hat{\mathcal{M}}^{(0)}=\hat{\mathcal{M}}$ are the initial inputs.
@@
$\hat{\mathcal{A}}^{(k)}=\mathbf{MHA}_{edge}(\hat{\mathcal{A}}^{(k-1)},\hat{\mathcal{L}}^{(k-1)},\hat{\mathcal{RPE}}^{(k-1)})$ (14)
@@ III-E Difficulty-guided decoder
To capture features of plausible future trajectories in the scene, we introduce a Difficulty-Guided Decoder to obtain the future trajectories of agents that are relatively easier to predict. Firstly, we decode the actor features $\hat{\mathcal{X}}^{(l)}$ in scence-centric, which have undergone multiple layers of feature interactions. Each agent receives six predicted trajectories. Following the ADAPT [4] trajectory decoder to get higher quality decoded trajectories by predicting endpoints for refinement. The decoding process can be expressed as follows:
^$\hat{\mathcal{E}}=\mathbf{MLP}_{\text{end}}(\hat{\mathcal{X}}^{(l)})+\mathbf{MLP}_{\text{refine}}(\text{cat}[\hat{\mathcal{X}}^{(l)},\hat{\mathcal{E}}])$ (15)
$\hat{\mathcal{P}}=\text{cat}[\mathbf{MLP}_{\text{traj}}(\text{cat}[\hat{\mathcal{X}}^{(l)},\hat{\mathcal{E}}]),\hat{\mathcal{E}}]$ (16)
where endpoints $\hat{\mathcal{E}}$ are predicted and refined by concatenating with the agent features $\hat{\mathcal{X}}^{(l)}$, then the trajectories $\hat{\mathcal{P}}$ are predicted and merged with the refined endpoints.
@@
0:  Predicted trajectories $\hat{\mathcal{P}}$, Predicted endpoints $\hat{\mathcal{E}}$, threshold $\tau$
0:  Masked trajectories $\hat{\mathcal{P}}_{masked}$
1:  $\hat{\mathcal{P}}_{masked}=set()$
2:  ¯^.$\bar{\mathcal{E}}\leftarrow\hat{\mathcal{E}}.mean()$
@@
4:     $\mathcal{D}=\|\hat{\mathcal{E}}-\bar{\mathcal{E}}\|_{2}$
5:     $\bar{\mathcal{D}}\leftarrow\mathcal{D}.mean()$
6:     if $\bar{\mathcal{D}}\leq\tau$ then
7:        $\hat{\mathcal{P}}_{masked}$.add($\hat{\mathcal{P}}_{l}$)
@@
9:  return  $\hat{\mathcal{P}}_{masked}$
The Argoverse 1 dataset show that the predicted trajectories of most of the vehicles in the scenarios exhibit a high degree of concentration, with more than 90% of the samples of straight ahead trajectories in each case [21]. This suggests that the motion patterns of most agents can be easily captured, reflecting real-world traffic scenarios. In order to prevent higher difficulty prediction trajectories from entering the subsequent modules and generating cumulative errors, we introduced a difficulty masker to filter the initial prediction results. ScenceTransformer[22], inspired by recent approaches to language modeling, has pioneered the idea of using a masking strategy as a query to the model, enabling it to invoke a single model to predict agent behavior in multiple ways (marginal or joint prediction). This mechanism in this paper is intended to mask out the initial prediction trajectories of the more difficult agents, thus providing better control over the prediction accuracy. This process can be formalized as Algo. 1.